OpenStudy (anonymous):

simplify: (√ 3+2) (√ 3-5√ 2)

OpenStudy (anonymous):

3-5sqrt6+2sqrt3-10sqrt2

OpenStudy (anonymous):

must be tired. lets try that again \[(\sqrt{3}+2)(\sqrt{3}-5\sqrt{2})\] \[\sqrt{3}\times \sqrt{3}-5\sqrt{6}+2\sqrt{3}-10\sqrt{2}\] \[3-5\sqrt{6}+2\sqrt{3}-10\sqrt{2}\]

OpenStudy (anonymous):

this, but the way, is called "multiplying" not "simplify" although i am sure the directions said "simplify"

OpenStudy (my4babiesandme):

wasnt sure what to do lol

OpenStudy (anonymous):

yea well basically ur supposed to distribute. use FOIL